OpenStudy (anonymous):

simplify: 4 over 2 + radical 3

OpenStudy (anonymous):

So 4/2=2. So 2+radical3 is simplified, since you can not do anything more with the radical 3

OpenStudy (anonymous):

\[4\div \left( 2 + \sqrt{3} \right)\]

OpenStudy (anonymous):

Is that what you meant?

OpenStudy (anonymous):

its 4 over 2+ radical 3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Or did you mean 4/(2+rad3)

OpenStudy (anonymous):

Ah nevermind, my computer loaded up what dashingblock typed earlier. :)

OpenStudy (anonymous):

Okay, so you can't have a radical in the denominator. If you multiply the radical by itself directly, then you get just the number under the radical. But you have to make sure whatever you're multiplying by is a one. Let me show you in mathematical terms. Dr. Pepper, shall I take this one and you can help another? :P

OpenStudy (anonymous):

So: 2 ________ 2 + rad3 Rad 3 times rad 3 gives you rad 9, right?

OpenStudy (anonymous):

\[\sqrt{9}\]And that is equal to the number three since the square root of 9 is 3. This means you now have: 2 _____ 2+3 So really, 2 ____ 5 But there's a problem because if you are going to multiply one part of a fraction (say, the denominator) by something, you must multiply the other part (numerator) by the same thing.

OpenStudy (anonymous):

So now its: 2 multiplied by \[\sqrt{3}\] __________________ 5

OpenStudy (anonymous):

Oh, dang, please replace all of my numerator 2 with numerator 4! Whoops, sorry! This is the correct answer : \[4\sqrt{3} \div 5\]